When 1.81 g of Vitamin K is dissolved in 25.0 g of camphor, the freezing point of the solution is lowered by 6.06 C. Look up the freezing paint and the Kf constant for camphor in the Colligative Constants table.

Calculate the molar mass of Vitamin K.
Answer #1

Melting point of camphor = 175 0C

Kf of camphor = 37.7 0C/m

Depression in freezing point = molality of solution * Kf

Solving, we get: molality = 6.06 / 37.7 = 0.161 m

Let molar mass of Vitamin K = x. Moles of Vit K = mass / molar mass = 1.81 / x moles

molality = moles of Vitamin k / mass of solution in kg = (1.81 / x) / (26.81 * 10-3) = 67.51 / x

Equating this to the molality we obtained above: 67.51 / x = 0.161

Solving, we have: x = 419.32 g/mol